OpenStudy (superhelp101):

What is a possible first step to eliminate the variable y in the following system? 4x + 6y = 1 5x - 3y = 8 Multiply the first equation by 5. Multiply the first equation by -2. Multiply the second equation by 2. Multiply the second equation by -3.

OpenStudy (superhelp101):

can someone plz help me? @ShailKumar

OpenStudy (superhelp101):

@genius12 @jim_thompson5910 @King.Liang @Luigi0210

OpenStudy (superhelp101):

@wio @w3ath3rfr3ak @Dark_Riddles @Dsal @ShailKumar @sweetburger do you think you can help me?

OpenStudy (superhelp101):

do u think you can help me on this problem @ShailKumar

OpenStudy (superhelp101):

:/ plz

OpenStudy (superhelp101):

could it be c? wait.. idk

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Multiply the second equation by 2.

OpenStudy (anonymous):

You'll get 2*(5x - 3y = 8) ==> 10x-6y = 16 Now you have two equations: 4x + 6y = 1 and 10x-6y = 16 Add them you get: 14 x = 17 See y is eliminated.

OpenStudy (superhelp101):

so it is C. Thank you so much!!! Can you help me with just one more? @ShailKumar

OpenStudy (anonymous):

Post that separately and tag me.

OpenStudy (superhelp101):

sure thxx